OpenStudy (britnee_paige99):

-1(5 + 3x)(5 + 3x) is this a perfect square trinomial?

Directrix (directrix):

Look at this.

Directrix (directrix):

-1(5 + 3x)(5 + 3x) = - ( 5 + 3x ) ^2

Directrix (directrix):

( 5 + 3x ) ^2 is a perfect square trinomial when multiplied and simplified. But that is not the question.

Directrix (directrix):

-( 5 + 3x ) ^2 is the question. Is that a perfect square trinomial.

Directrix (directrix):

-( 5 + 3x ) ^2 = -9x^2 - 30x -25

Directrix (directrix):

Is the first term a positive perfect square?

rebeccaxhawaii (rebeccaxhawaii):

^^

OpenStudy (skullpatrol):

$$-9x^2 =?=a^2$$ $$-(3x)^2$$